OpenStudy (anonymous):

Solve using linear equations: Pierre has found that his new car gets 33 miles per gallon on the highway and 27 miles per gallon in the city. He recently drove 317 miles on 10 gallons of gasoline. How many miles did he drive on the highway? How many miles did he drive in the city?

OpenStudy (wolf1728):

Highway driving 'h' gets 33 mpg City driving 'c' gets 27 mpg Let 'h' be gallons for highway and 'c' be gallons used in city We can make 2 equations A) 33h + 27c = 317 B) h + c = 10 Multiplying B by -33 -33h -33c = -330 Adding equation A to the above yields -6c = -13 c = 2.1666666667 gallons h = 10 -c = 10-2.1666666667 = 7.8333333333 gallons The question asked how much of this is city driving and highway driving. 7.8333333333 * 33 = 258.5 miles of highway driving 2.1666666667 * 27 = 58.5 miles of city driving 258.5 + 58.5 = 317 miles

OpenStudy (wolf1728):

Thanks for the medal :-)

OpenStudy (anonymous):

Thank YOU for the answer, it helped a lot :)